# Interpret the Internal States of Recommendation Model with Sparse Autoencoders

### v2

@@ 3.2.2. Module Architecture
Unlike standard autoencoders, RecSAE scales the input by a factor of $s$ relative to the original activation dimension and incorporates an additional sparsity objective to promote mono-semanticity (Templeton et al., 2024) in the latent space. A vanilla approach to achieve sparsity is to apply an $L_{1}$ regularization term to the training loss. However, this often results in a dense distribution of low-activation latents, which does not align to achieve $L_{0}$ sparsity in the latent space $S_{Latent}$.
To address this, RecSAE employs a Top-k activation layer, which explicitly controls the number of active latents after encoding. As illustrated by the equations below, $S_{Latent}$ denotes the sparse latent vectors with $k$ latents activated and the others set to 0, and $\hat{S}_{Internal}$ is the reconstructed activations. $M_{Pe}$ and $M_{Pd}$ represent the encoder module and decoder module of the RecSAE module. $W_{enc}$, $b_{pre}$, $W_{dec}$ represent the encoder weight, bias and decoder weight. RecSAE hyper-parameters, scale size $s$ and sparsity factor $k$ are discussed in Section 5.3.
@@ 3.2.3. Training Loss
RecSAE adopts an unsupervised learning approach, minimizing the $L_{2}$ distance between $S_{Internal}$ and $\hat{S}_{Internal}$ to train the module. This process ensures that the large, sparse latent space $S_{Latent}$ effectively encodes the information from the internal activations. Thus, interpretations of the latent space reflect information encoded in the interpreted model.
@@
In addition to the primary optimization objective, training RecSAE with Top-k activation introduces a key challenge: preventing ”dead” latents, units that remain inactive during training or inference, to fully utilize the expanded latent space. While activating $k$ latents per input improves efficiency and interpretability, it also means that inactive latents receive no gradient updates from the reconstruction loss and therefore are not effectively trained. To address this, we introduce an auxiliary loss, drawing inspiration from previous work (Gao et al., 2024; Jermyn and Templeton, 2024). The auxiliary loss targets these inactive latents during training, guiding their learning even though they are not involved in the main reconstruction process. Formally, given the reconstruction error $e$, the auxiliary loss $\mathcal{L}_{aux}$ is defined as the $L_{2}$ distance between the main reconstruction error and the reconstruction using top $k_{auxD}$ dead latents, as shown in the equations below. Since the encoder preceding the Top-k activation layer is shared, adding this auxiliary loss incurs limited computational overhead. The training loss is the linear combination of $\mathcal{L}$ and $\mathcal{L}_{aux}$. The effectiveness of RecSAE training designs is validated in Section 5.4.
(2) $\begin{split}\text{loss}&=\mathcal{L}+\alpha\mathcal{L}_{aux}\\ \mathcal{L}&=||S_{Internal}-\hat{S}_{Internal}||^{2}_{2}\\ \mathcal{L}_{aux}&=||e-\hat{e}||^{2}_{2},\quad\text{where }e=S_{Internal}-\hat{S}_{Internal},\\ \hat{e}&=W_{dec}S_{L}^{\prime},\ S_{L}^{\prime}=\text{Top-k}_{auxD}(W_{enc}(S_{Internal}-b_{pre})).\end{split}$
@@
With the above structure and training strategy, RecSAE is capable of probing various types of recommendation models. While applied to well-trained models, RecSAE allows for interpretation without altering model functions. Moreover, by modifying latent activations and replacing the original activations with their modified reconstructions, RecSAE facilitates targeted de-biasing of models. This potential is discussed in Section 5.5.
3.3. RecSAE Concept Construction
@@
After training, RecSAE establishes the mapping relationships between inputs, RecSAE latent activations, and model outputs. Using these links, we design an automated, closed-loop process for generating and validating interpretations for each latent.
@@
RecSAE uses the test interactions for concept construction and verification. As shown in the lower part of Figure  2, it first records the input, latent activations, and recommendation outputs during base model inference on the test set. Next, it divides the continuous, unbounded activation values in the latent space into 10 intensity levels (ranging from 1 to 10), with no activation corresponding to level 0. Then, RecSAE selects the top 2N most activated cases for each latent, randomly assigning N cases for concept construction and reserving the remaining N for verification. Additionally, N non-activated cases for each latent are randomly selected for verification. In line with OpenAI’s setting to interpret neurons with language models (Bills et al., 2023), we set N=5 in this study.
@@
For concept construction, RecSAE generates textual descriptions for each latent, transforming latents into textual concepts, as detailed in Subsection 3.3.1. For concept verification, RecSAE evaluates the interpretations by predicting activation levels on new inputs and calculating precision and recall scores to determine interpretation confidence, as outlined in Subsection 3.3.2.
3.3.1. Concept Description Construction
For latent interpretations, we analyze the highly activated cases. RecSAE incorporates item meta information (title and category if available) to form textual sequences for semantic concept constructions. Each sequence consists of user history interacted items, corresponding model prediction item and its activation level on the interpreted latent.
@@ 3.3.1. Concept Description Construction
As shown in Table 5, RecSAE latents effectively extract items with notable similarity. Conseuently, summarizing these similarities into coherent descriptions is a relatively straightforward task. Therefore, instead of relying on paired APIs with state-of-the-art large language models, we opt for a locally deployable model, Llama-3-8B-Instruct (AI@Meta, 2024), to generate concept descriptions in this paper. The instruction starts with a system prompt outlining the latent interpretation task, followed by a one-shot example that includes a query and the corresponding assistant response of concept descriptions. Then, the interpreted latents’ activation cases are provided for interpretation. The complete instruction with one-shot example for each dataset is presented in our anonymous repository.
3.3.2. Concept Description Verification
@@
With the automatically generated concepts, RecSAE further evaluates their quality through confidence scores. To this end, RecSAE uses concept descriptions to predict latent activations on new instances. Specifically, we use two types of test cases: highly activated cases (distinct from those used for concept constructions), and randomly selected cases with no activation. These two sets allow us to compute recall and precision, which together form the basis of the confidence scores.
@@
The verification process begins with a system prompt that outlines the activation-level prediction task, followed by a one-shot example. This example includes a user query containing the concept description, a test case, and the assistant’s predicted activation level. The template is available in the codebase. After predicting latent activation levels, RecSAE computes a confidence score for each concept. A test case is labeled as positive if its predicted activation exceeds 5; otherwise, it is labeled negative. The confidence score is computed as the proportion of correctly classified cases (i.e., true positives and true negatives) among tests, reflecting the accuracy of the concept description in representing latent semantics. In Section 5.2.3, human experts annotate concept quality with confidence scores $\geq$ 0.8 to further validate the effectiveness of RecSAE’s automatic concept construction and validation pipeline.
@@ 4. Experimental Setup
We implement the RecSAE framework based on a publicly available recommendation framework, ReChorus (Li et al., 2024), which is recommended by the RecSys Conference [^2] https://github.com/ACMRecSys/recsys-evaluation-frameworks.
@@ 4.1. Dataset
We utilize three widely used datasets: (1) Amazon (Leskovec et al., 2007): the Grocery & Gourmet Food subset of a large-scale Amazon Reviews dataset. (2) MovieLens 1M (Harper and Konstan, 2015): a rating data set from the MovieLens website. (3) LastFM (Schedl et al., 2022): listening records of the music platform Last.fm.
We use the built-in data processing in ReChorus: retain positive interactions, 5-core filter, and leave-one-out train, dev, and test split. Table 2 lists the statistics of three datasets. Both datasets contain ID-based interactions and item textual descriptions.
@@
For SAE training, we infer the interpreted recommendation model over the training set and record their internal activations at the probing position during model inference. With the same approach, we construct the valid and test set for RecSAE.
@@ 4.3. Baseline
RecSAE is a model-level interpretation framework that is not restricted to specific architectures. As shown in Table 1, the only comparable generalized model-level baseline is ECF (Du et al., 2023), a clustering-based method that extracts taste clusters from interactions and item tags. To further strengthen our comparison, we include the distillation method DistillEmb (Zhang et al., 2020b) as an additional baseline, though it is limited to jointly trained models. We also probe internal activations of base models for interpretation comparisons in Table 3. These 3 baselines allow us to evaluate RecSAE’s effectiveness in performance and interpretability across diverse methods.
@@ 4.4. Evaluation Metrics
RecSAE aims to interpret recommendation models, requiring that its latent activations reflect the internal states of the interpreted model. Thus, we evaluate both the reconstruction and interpretation performance of RecSAE latents.
For reconstruction, we evaluate the downstream recommendation performance by hit rate and NDCG at positions of 10 and 20 in test-all settings when the original activations are replaced with the reconstructions in the interpreted model, though this replacement is not needed during interpretation.
For interpretation, we refer to related work (Du et al., 2023) that measures the latent utilization through Silhouettes (Rousseeuw, 1987) (a measure of how similar an object is to its own cluster) and the number of interpreted concepts within different confidence score ranges. Specifically, Silhouettes evaluates intra- and inter-concept similarity. The similarity is measured using the cosine distance between item embeddings in the interpreted models. Intra-concept similarity quantifies the average distances between the two most activated items within each concept. Meanwhile, inter-concept similarity measures the average distances between pairs of distinct concepts. We calculate the above measures in concepts with a confidence score of 0.8 or higher. High intra- and lower inter-concept similarity indicate that the interpreted concepts are well-defined and distinct.
@@ 4.5. Hyper-parameter
For recommendation models, we use the tuned settings in the ReChorus framework. The embedding size for recommenders is 64, so sparsity $k$ in RecSAE should be less than 64. We perform hyperparameter tuning over the scale size {8, 16, 32}, the sparsity k {8, 16}, learning rate {1e-5, 5e-5, 1e-4} and batch size {4, 8, 16}. The impact of scale size and sparsity k is presented in Section 5.3. For final experiments, we set the scale size to 16 (leading to 1024 latents) and k to 8 for reconstruction and interpretation balance. The detailed hyper-parameter settings for interpreting each model are reported in run.sh within the codebase.
Table 3. RecSAE’s Overall Performance. We evaluate RecSAE’s training with downstream performance when replacing original activations with reconstructions. Note that this replacement is not required during interpretation. For interpretation performance, we measure the item similarity within and between concepts and evaluate interpreted concept numbers (All, confidence score = 1, $\geq$0.9, $\geq$0.8). The best metric on each dataset is bold and the second best is underlined.
Dataset Rec Model Interpret Method Reconstruction Interpretation
Downstream Performance ↑ Similarity #Concepts ↑

### v3

@@ 3.2.2. Module Architecture
Unlike standard autoencoders, RecSAE scales the input by a factor of $s$ relative to the original representation dimension and incorporates an additional sparsity objective to promote mono-semanticity (Templeton et al., 2024) in the latent space. A vanilla approach to achieve sparsity is to apply an $L_{1}$ regularization term to the training loss. However, $L_{1}$ is an imperfect approximation of $L_{0}$ and it introduces a bias of shrinking all positive activations toward zero (Gao et al., 2024).
To address this, RecSAE employs a Top-k activation layer, which explicitly controls the number of active latents after encoding. The Top-K activation layer has been proved to outperforms other activation layers (ReLU, ProLU STE, Gated) on sparsity-reconstruction frontier in sparse autoencoders (Gao et al., 2024). By contrast, using a sparsity‐inducing penalty term requires careful tuning of a regularization weight, which can vary across different recommendation types. Consequently, we adopt the Top-K activation mechanism to ensure stable sparsity control in RecSAE. The formal module structure is illustrated by the following equations, where $S_{Latent}$ denotes the sparse latent vectors with $k$ latents activated and the others set to 0, and $\hat{S}_{Internal}$ is the reconstructed representations. $M_{Pe}$ and $M_{Pd}$ represent the encoder module and decoder module of the RecSAE module. $W_{enc}$, $b_{pre}$, $W_{dec}$ represent the encoder weight, bias and decoder weight. The sensitivity of RecSAE to its hyper-parameters, scale size $s$ and sparsity factor $k$, is further discussed in Section 7.
@@ 3.2.3. Training Loss
RecSAE adopts an unsupervised learning approach, minimizing the $L_{2}$ distance between $S_{Internal}$ and $\hat{S}_{Internal}$ to train the module. This process ensures that the large, sparse latent space $S_{Latent}$ effectively encodes the information from the internal representations. Thus, interpretations of the latent space reflect information encoded in the interpreted model.
@@
In addition to the primary optimization objective, training RecSAE with Top-k activation introduces a key challenge: preventing ”dead” latents, units that remain inactive during training or inference, to fully utilize the expanded latent space. While activating $k$ latents per input improves efficiency and interpretability, it also means that inactive latents receive no gradient updates from the reconstruction loss and therefore are not effectively trained. To address this, we introduce an auxiliary loss, drawing inspiration from previous work (Gao et al., 2024; Jermyn and Templeton, 2024). The auxiliary loss targets these inactive latents during training, guiding their learning even though they are not involved in the main reconstruction process. Formally, given the reconstruction error $e$, the auxiliary loss $\mathcal{L}_{aux}$ is defined as the $L_{2}$ distance between the main reconstruction error and the reconstruction using top $k_{auxD}$ dead latents, as shown in the equations below. Since the encoder preceding the Top-k activation layer is shared, adding this auxiliary loss incurs limited computational overhead. The training loss is the linear combination of $\mathcal{L}$ and $\mathcal{L}_{aux}$.
[…]
@@
With the above structure and training strategy, RecSAE is capable of probing various types of recommendation models. While applied to well-trained models, RecSAE allows for interpretation without altering model functions. Moreover, by modifying latent activations and replacing the original representations with their modified reconstructions, RecSAE facilitates targeted tuning of models. This potential is discussed in Section 8.
3.3. RecSAE Concept Construction and Verification
@@
After training, RecSAE establishes the mapping relationships between inputs, RecSAE latent activations, and model outputs. Using these links, we use an automated, closed-loop pipeline for generating and validating semantic concepts for each latent.
@@
For interpretation dataset construction, we build an interpretation training set for concept construction and an interpretation test set for concept verification. The recommendation test set is reserved for interpretation verification, and the recommendation training set is used for interpretation construction. This setup ensures that RecSAE concept construction is conducted without data leakage. As illustrated in the lower part of Figure 1, the process first records the input, latent activations, and recommendation outputs obtained during the base recommendation model’s inference on both the training and test sets. Note that for the main experiment results, the recommendation outputs refer to the top-1 recommended item. Next, we partition the continuous latent activations into 11 intensity levels (0–10): level 0 denotes non-activation, and levels 1–10 are equal-frequency bins derived from the training-set distribution. Then, RecSAE selects the top 2N most activated cases for each latent, randomly assigning N cases for concept construction and reserving the remaining N for verification. Additionally, N non-activated cases for each latent are randomly selected for verification. In line with OpenAI’s setting to interpret neurons with language models (Bills et al., 2023), we set N=5 in this study.
@@
For concept construction, RecSAE generates textual descriptions for each latent, transforming latents into textual concepts, as detailed in Subsection 3.3.1. For concept verification, RecSAE evaluates the interpretations by predicting activation levels on new inputs and calculating precision and recall scores to determine interpretation confidence, as outlined in Subsection 3.3.2. We employ two methods in the above two-stage pipeline: (1) the TF-IDF (term frequency–inverse document frequency) algorithm for term-level statistical summarization; (2) a large language model for semantic-level analysis. The statistical method provides a reliable baseline, while the LLM offers a strong reference point for the upper bound of automatic interpretation performance.
3.3.1. Construction
@@
RecSAE can latents effectively extract items with notable similarity, as further shown in Table 5. Consequently, summarizing these similarities into coherent descriptions is a relatively straightforward task. Therefore, we adopt both term-level algorithm and an open-sourced LLM, instead of relying on paid APIs with state-of-the-art close-sourced large language models. We opt the TF-IDF algorithm (Christian et al., 2016) and a locally deployable model, Llama-3-8B-Instruct (AI@Meta, 2024), to generate concept descriptions in this paper. For the TF-IDF algorithm, we construct the textual representation for each latent using the metadata of items in the interpretation set, and identify its concept as the term with the highest TF-IDF score.
For LLM interpretation, the instruction starts with a system prompt outlining the latent interpretation task, followed by a one-shot example that includes a query and the corresponding assistant response of concept descriptions. Then, the interpreted latents’ activation cases are provided for interpretation. The complete instruction with one-shot example for each dataset is presented in our anonymous repository.
Instruction for Concept Construction ¡Role.SYSTEM¿
We’re studying neurons in a recommendation model that is used to recommend {item_category}. Each neuron looks for some particular concepts in {item_category}. Look at the parts of the {item_name} the neuron activates for and summarize in a single sentence what the neuron is looking for. Don’t list examples of words.
The activation format is {item_name}¡tab¿activation. Activation values range from 0 to 10. A neuron finding what it’s looking for is represented by a non-zero activation value. The higher the activation value, the stronger the match.
¡Role.USER¿
Neuron 0
Activations:
{example_item_activation_sequences}
Explanation of neuron 0 behavior: the main thing this neuron does is find
¡Role.ASSISTANT¿
{example_explanation}
¡Role.USER¿
Neuron {latent_id}
Activations:
{item_activation_sequences}
Explanation of neuron {latent_id} behavior: the main thing this neuron does is find
Specifically, the above instruction analyzes item activation sequences. Each sequence includes the user’s interaction history, the model’s predicted item, and the corresponding activation level on the interpreted latent. An example of such a sequence is shown below:
¡start¿
Bob’s Red Mill Organic Scottish Oatmeal, 20-Ounce Bags (Pack of 4) (0)
Funky Monkey Snacks Carnaval Mix, Freeze-Dried Fruit, 1-Ounce Bags (Pack of 12) (0)
Mrs. May’s Trio Bar Variety Pack, 1.2-Ounce bars (Pack of 20) (0)
Mrs. May’s Strawberry Trio Bar, 1.2-Ounce Bars (Pack of 24) (10)
¡end¿
Since the model activates the relevant latent when predicting the recommended item, the activations occur on the predicted item. As shown in the example, a latent activation sequence consists of the user’s historical interacted items, serving as contextual input with zero activation on the latent, and the model’s predicted items, which exhibit the corresponding activation levels on that specific latent. With the above instruction and the activated sequence information of each latent, the pipeline use the Llama model to form the interpreted concepts.
3.3.2. Verification
@@
To assess the quality of the automatically generated concepts, RecSAE evaluates them through confidence scores. Specifically, it leverages the concept to predict latent activations on new instances. Specifically, we use two types of test cases: highly activated cases (distinct from those used for concept constructions), and randomly selected cases with no activation. If a concept appears in a negative case, it is counted as a failure; if it appears in a positive case, it is counted as a hit. The use of both sets enables the calculation of confidence scores and ensures that the verification process is discriminative.
For the TF-IDF method, verification is performed through exact term matching. If the extracted concept term appears in the metadata of an item within the verification set, it is considered a hit; otherwise, it is marked as a miss. The interpretation confidence score for each latent is then computed as the proportion of hit cases among all cases in the verification set.
@@
For LLM method, the verification process begins with a system prompt that outlines the activation-level prediction task, followed by a one-shot example. This example includes a user query containing the concept description, a test case, and the assistant’s predicted activation level. The template is available in the codebase. After predicting latent activation levels, RecSAE computes a confidence score for each concept. A test case is labeled as positive if its predicted activation exceeds 5; otherwise, it is labeled negative. The confidence score is computed as the proportion of correctly classified cases (i.e., true positives and true negatives) among tests, reflecting the accuracy of the concept description in representing latent semantics. The formal calculation of the confidence score is given in Equation 4, where pos and neg denote the sets of activated and non-activated test cases, respectively; $level(x)$ is the LLM-predicted activation level for test case $x$; thresthres is the classification threshold; and $\mathbbm{1}$ is the indicator function.
(4) $Precision_{concept}=\frac{\sum_{p\in\text{pos}}\mathbbm{1}[level(p)>thres]+\sum_{n\in\text{neg}}\mathbbm{1}[level(n)<=thres]}{|\text{pos}|+|\text{neg}|}$
Instruction for Concept Verification ¡Role.SYSTEM: ’system’¿ ”We’re studying neurons in a recommendation model that is used to recommend {item_category}. Each neuron looks for some particular concepts in {item_category}. Look at an explanation of what the neuron does, and try to predict its activations on a particular {item_name}. The activation format is {item_name}¡tab¿activation, and activations range from 0 to 10. Most activations will be 0. Now, we’re going predict the activation of a new neuron on a single {item_name}, following the same rules as the examples above.
¡Role.USER: ’user’¿ Neuron {example_latent_id}
Explanation of neuron {example_latent_id} behavior: the main thing this neuron does is find {example_explanation}
Sequence:
{example_sequence}
Last product in the sequence:
{example_last_item}
Last product activation, considering the product in the context in which it appeared in the sequence:
¡Role.ASSISTANT: ’assistant’¿ {example_last_item_activation}
¡Role.USER: ’user’¿ Neuron {test_latent_id}
Explanation of neuron {test_latent_id} behavior: the main thing this neuron does is find {test_explanation}
Sequence:
{test_sequence}
Last product in the sequence:
{test_last_item}
Last product activation, considering the product in the context in which it appeared in the sequence:
In Section 5.4, we further validate the effectiveness of RecSAE’s automated concept construction and verification pipeline through expert annotation. Concepts with confidence scores $\geq 0.8$ are selected for human evaluation to assess their quality.
@@ 4. Experimental Setup
In this section, we describe the experimental setup used in our study. We implement the RecSAE framework based on a publicly available recommendation framework, ReChorus (Li et al., 2024), which is recommended by the RecSys Conference [^2] https://github.com/ACMRecSys/recsys-evaluation-frameworks. Section 4.1 provides an overview of the four publicly available and widely used datasets employed in our experiments. Section 4.2 introduces the three types of recommendation models used for interpretations. Section 4.3 outlines the baseline models for comparison with RecSAE in terms of interpretation. Section 4.4 details the evaluation metrics used to assess both interpretation and RecSAE reconstruction performance. Section 4.5 presents the hyper-parameters used in the RecSAE implementation, and their sensitivity experiments is further discussed in experiment Section 7.
@@ 4.1. Dataset
We utilize four widely used datasets: (1) Amazon (Leskovec et al., 2007): the Grocery & Gourmet Food subset of a large-scale Amazon Reviews dataset. (2) MovieLens 1M (Harper and Konstan, 2015): a rating data set from the MovieLens website. (3) LastFM (Schedl et al., 2022): listening records of the music platform Last.fm. (4) Yelp (Yelp, 2018): a dataset of Yelp businesses, reviews, and user data from Yelp Challenge 2018.
We adopt the built-in data settings and processing procedures provided by ReChorus. Specifically, for the Amazon dataset, positive interactions refer to users who have reviewed items; for the MovieLens dataset, positive interactions correspond to ratings equal to or above 4; for the LastFM dataset, positive interactions represent listening events between a user and a track; for the Yelp dataset, positive interactions represent reviews with ratings equal to or above 4. For data processing, the framework applies a 5-core filter, and performs a leave-one-out split for training, validation, and testing sets. Table 2 lists the statistics of four datasets. Both datasets contain ID-based interactions and item textual descriptions.
@@
Yelp 180,156 79,601 2,559,541
For SAE training, we infer the interpreted recommendation model over the training set and record their internal representations at the probing position during model inference. With the same approach, we construct the valid and test set for RecSAE.
@@ 4.3. Baseline
RecSAE is a model-level interpretation framework that is not restricted to specific architectures. To the best of our knowledge, no existing work is applicable to interpreting general, graph-based, and sequential recommendation models in a unified manner. Therefore, for comparison purposes, we probe the internal representations of base models for interpretation comparisons in Table 3 and 4. These baselines allow us to evaluate RecSAE’s effectiveness in performance and interpretability across diverse methods.
@@ 4.4. Evaluation Metrics
RecSAE aims to interpret recommendation models, requiring that its latent reflect the internal states of the interpreted model. Thus, we evaluate both the reconstruction and interpretation performance of RecSAE latents.
For reconstruction, we evaluate the downstream recommendation performance by hit rate and NDCG at positions of 10 and 20 in test-all settings when the original representations are replaced with the reconstructions in the interpreted model, though this replacement is not needed during interpretation. The reconstruction results are shown in Section 6.
For interpretation, we refer to related work (Du et al., 2023) that measures the latent utilization through Silhouettes (Rousseeuw, 1987) (a measure of how similar an object is to its own cluster) and the number of interpreted concepts within different confidence score ranges. Specifically, Silhouettes evaluates intra- and inter-concept similarity. The similarity is measured using the cosine distance between item embeddings in the interpreted models. Intra-concept similarity quantifies the average distances between the two most activated items within each concept. In addition, inter-concept similarity measures the average distances between pairs of distinct concepts. We calculate the above measures in concepts with a confidence score of 0.8 or higher. High intra- and lower inter-concept similarity indicate that the interpreted concepts are well-defined and distinct. The interpretation results are shown in Section 5.
@@ 4.5. Hyper-parameter
For recommendation models, we use the tuned settings in the ReChorus framework. The embedding size for recommenders is 64, so sparsity $k$ in RecSAE should be less than 64. We perform hyperparameter tuning over the scale size {8, 16, 32}, the sparsity k {8, 16}, learning rate {1e-5, 5e-5, 1e-4} and batch size {4, 8, 16}. The impact of scale size and sparsity k is presented in Section 7. For final experiments, we set the scale size to 16 (leading to 1024 latents) and k to 8 for reconstruction and interpretation balance. The detailed hyper-parameter settings for interpreting each model are reported in run.sh within the codebase.
In the following experiment sections, we present the interpretation performance, reconstruction performance, parameter sensitivity experiments, and targeted tuning.
5. Interpretation Performance
RecSAE consists of two phases: the unsupervised learning phase, which forms the interpretable latents, and the interpretation phase, which generates semantic concepts based on the items that activate each latent. Thus, we evaluate both the quality of the learned latents and the interpreted concepts.
In the following sections, we: 1) Assess the quality of the latents learned by RecSAE; 2) Present the interpreted concepts, focusing on their quantity, activation patterns, concrete examples, and concept survey; 3) Conduct a human evaluation to assess the validity and clarity of the interpreted concepts produced by RecSAE.
5.1. RecSAE Extracted Latents
Table 3. Evaluation of RecSAE Latents: Intra- and Inter-Latent Similarity for RecSAE across Interpreted Models and Datasets. The best performance in each setting is highlighted in bold. Higher intra-latent similarity and lower inter-latent similarity show that RecSAE Learns valid and distinct latents.
Rec Model Method Amazon MovieLens LastFM Yelp
Intra ↑ Inter ↓ Intra ↑ Inter ↓ Intra ↑ Inter ↓ Intra ↑ Inter ↓
Base 0.1764 0.2363 0.1804 0.4740 0.1437 0.1495 0.5161 0.6492
BPRMF RecSAE 0.3056 0.1663 0.3795 0.2998 0.3058 0.1431 0.3594 0.2626
Improve +73.30% -29.65% +110.33% -36.75% +112.85% -4.24% -30.37% -59.55%
Base 0.2161 0.2463 0.4430 0.2938 0.1818 0.1565 0.3433 0.3235
LightGCN RecSAE 0.4540 0.1496 0.6225 0.2704 0.4672 0.1365 0.6389 0.2248
Improve +110.03% -39.23% +40.54% -7.97% +156.95% -12.79% +86.08% -30.50%
Base 0.3777 0.1808 0.4452 0.2071 0.2652 0.1755 0.1853 0.1263
SASRec RecSAE 0.7147 0.1587 0.6980 0.1745 0.5114 0.1579 0.5836 0.1307
Improve +89.21% -12.20% +56.80% -15.76% +92.81% -10.03% +215.00% +3.53%
Base 0.3418 0.1405 0.5119 0.2303 0.2332 0.1626 0.3282 0.1261
TiMiRec RecSAE 0.6288 0.1366 0.6564 0.1895 0.5332 0.1470 0.6635 0.1375
Improve +83.96% -2.77% +28.22% -17.71% +128.60% -9.56% +102.18% +9.00%
We begin our evaluation of RecSAE by analyzing the quality of extracted latents after the unsupervised learning phase. Specifically, we aim to assess two essential properties of the learned latents: (1) Intra-latent coherence, indicating that items highly activated by the same latent are similar to one another; and (2) Inter-latent distinctiveness, meaning that items activated by different latents are meaningfully dissimilar. To quantify these properties, we compute intra-latent similarity and inter-latent similarity metrics across different datasets and recommendation models. Higher intra-latent similarity suggests tighter clustering of semantically similar items under a shared latent, while lower inter-latent similarity implies that different latents capture non-overlapping aspects of item semantics.
Table 3 reports the results for four representative recommendation models (BPRMF, LightGCN, SASRec, and TiMiRec) across four benchmark datasets (Amazon Grocery, MovieLens, LastFM and Yelp). For each model and dataset, we compare the base representation (i.e., the original model’s internal encoding) with the RecSAE learned sparse latents.
RecSAE significantly improves intra-latent similarity, confirming that items grouped under the same latent activation are more semantically cohesive. For instance, in the Amazon dataset, the intra-latent similarity of BPRMF increases from 0.1764 to 0.3056 (a +73.30% improvement). Simultaneously, RecSAE reduces inter-latent similarity, ensuring greater separation between different latents. In the same setting, the inter-latent similarity for BPRMF drops from 0.2363 to 0.1663 (-29.65%). This pattern of improved intra-latent cohesion and reduced inter-latent overlap holds consistently across all evaluated settings, including both general (e.g., BPRMF), graph-based (LightGCN), and sequential (e.g., SASRec, TiMiRec) models. There is one special case on BPRMF trained on Yelp, where both intra- and inter-latent similarities are generally lower. Note that these similarity metrics represent only one aspect of evaluating RecSAE; the primary utility of RecSAE lies in producing interpretable concepts and enabling subsequent targeted tuning.
These results highlight RecSAE’s effectiveness in structuring the interpretable latent space purely from the internal representations of recommendation models through unsupervised learning. By consolidating semantically related items and disentangling unrelated patterns, RecSAE achieves two essential properties that underpin mono-semantic latents in recommendation systems.
Finding 1: RecSAE learns valid and distinct latents.
5.2. RecSAE Interpreted Concepts
After extracting interpretable latents from the internal representations of recommendation models, we further transform these latents into human-readable semantic concepts. In this section, we present analysis of this interpretation process, including quantitative overview of interpretation performance, latent activation situations, case study of representative concepts, and broader concept surveys that highlight the effectiveness, coverage, and emerging patterns of interpretation.
5.2.1. Overall Performance